Subject: REQUIRED RATE OF RETURN Assume that the risk-free rate is 7% and the required return on the market is 11%. What is the required rate of return on a stock with a beta of 1.7? Round your answer to two decimal places.
Answer #1

Rm=11.1. RF = 7%. B = 1-7 Required return = Rp + (Rm - RfJX B. = 7+ (11 - 7) x 1• 7 = 13.80%